### v2

# INRetouch: Context Aware Implicit Neural Representation for Photography Retouching

Abstract
Professional photo editing remains challenging, requiring extensive knowledge of imaging pipelines and significant expertise. With the ubiquity of smartphone photography, there is an increasing demand for accessible yet sophisticated image editing solutions. While recent deep learning approaches, particularly style transfer methods, have attempted to automate this process, they often struggle with output fidelity, editing control, and complex retouching capabilities. We propose a novel retouch transfer approach that learns from professional edits through before-after image pairs, enabling precise replication of complex editing operations. To facilitate this research direction, we introduce a comprehensive Photo Retouching Dataset comprising 100,000 high-quality images edited using over 170 professional Adobe Lightroom presets. We develop a context-aware Implicit Neural Representation that learns to apply edits adaptively based on image content and context, requiring no pretraining and capable of learning from a single example. Our method extracts implicit transformations from reference edits and adaptively applies them to new images. Through extensive evaluation, we demonstrate that our approach not only surpasses existing methods in photo retouching but also enhances performance in related image reconstruction tasks like Gamut Mapping and Raw Reconstruction. By bridging the gap between professional editing capabilities and automated solutions, our work presents a significant step toward making sophisticated photo editing more accessible while maintaining high-fidelity results.
Figure 1: Comparison of different approaches for style transfer in the context of reference based image editing.
1 Introduction
Photos are an integral part of our lives, used for sharing information, expressing experiences, showcasing art, and storytelling. This widespread usage drives a demand among all types of photographers for increasingly sophisticated photo editing tools. These tools allow users to retouch their images to achieve accurate or artistic representations of captured scenes. Professional photographers often rely on advanced software like Adobe Lightroom [1] and PhotoLab [2] to edit RAW images, requiring a strong grasp of image processing concepts such as contrast, saturation, white balance, and tone mapping. In contrast, smartphone users frequently apply presets and filters, which are typically built on predefined Look-Up Tables (LUTs) for basic, global adjustments [11, 12], with unfortunately very limited options.
With the rise of learning-based methods, new approaches to image manipulation have emerged. Techniques such as style transfer allow users to specify a reference style image, which a neural network then applies to an input image [15, 20, 19]. Examples of this process are shown in Figure 1. However, these methods are mainly suited for artistic applications, such as transferring the style of a painting onto a photograph, rather than high-fidelity photography editing, where subtle adjustments that retain the realism of the original scene are desired [12].
For more photo-realistic applications, recent advancements have been made in photo-realistic style transfer[32, 28]. Another group of works approaches the problem as a deterministic color mapping[19, 23], also known as color style transfer. These methods are widely used in the industry due to their ability to avoid output artifacts and produce accurate results. Nevertheless, they often lack flexibility, primarily allowing only global adjustments and failing to provide more refined modifications, such as local adjustments (e.g., enhancing just the sky in an image) or artistic effects like vignetting. This observation prompts us to consider the following questions:
Why previous methods are not enough?
Previous works [15, 32, 19] rely on a reference image to define the target style, leaving the network to determine which elements of the input image should change to match that style. This approach provides no direct control over the specific alterations applied to the input and often results in unintended content changes, particularly when the reference image has different content. While deterministic color transfer methods [19, 23] yield more reliable results with fewer distortions, they are largely restricted to global and subtle color adjustments. These works lack the flexibility to apply other popular modifications, such as introducing artistic noise, making localized adjustments, or adding vignetting effects.
How can we overcome previous limitations?
Drawing inspiration from the concept of image analogies [18], we propose a novel approach for automatic photo editing by learning from examples. By supplying the model with pairs of before-and-after edited images, it provides the opportunity to learn the specific edits applied and replicate them on new input images. This approach frames the task as a deterministic retouching transfer—extending beyond basic color and general appearance adjustments. Additionally, it allows precise control over changes, as the model transfers only the differences present in the reference example.
To promote such a research direction, we created a unique photo retouching dataset (PRD) using over 170 Adobe Lightroom presets crafted by professional photographers. These presets were applied across images with diverse scenes, lighting conditions, and subjects, resulting in a dataset of approximately 100,000 high-quality retouched images. Unlike prior works [19, 23], our dataset includes complex global and local transformations, not limited to simple color adjustments. To our knowledge, this is the most realistic and challenging dataset of its kind.
How to learn the edits ?
To achieve this goal, we propose a novel Retouch Transfer method leveraging Implicit Neural Representation (INR). INRs offer a powerful approach for compressing data into compact forms and interpolating missing information [42, 10]. We harness this capability by reframing the image editing task as adaptive color transformation. Our approach introduces a unique INR architecture that incorporates spatial and contextual awareness, enabling complex, localized, and adaptive edits. Unlike traditional pipelines [19, 23], our method requires no pretraining and can learn edits from just a single example, allowing it to handle editing inconsistencies within the dataset effectively. Additionally, we demonstrate the effectiveness of our INR architecture in other image restoration tasks, such as Gamut Mapping [26] and Raw Reconstruction [27], enhancing performance over conventional INR architectures with minimal computational and memory costs.
To summarize, we study the possibility of achieving the retouch transfer, from a pre-edited reference image. This new research direction offers greater control over the applied edits, closely resembling the professional photo editing process. This flexibility makes it a powerful tool for transferring edits from previously edited images, minimizing the need for specialized editing software (e.g., Lightroom for presets) and enabling edit transfer from online images. Additionally, the adaptability and efficiency of our method allow it to surpass limited 3D LUT color filters, enabling the creation of more complex image style transformations. This also positions our method as a strong alternative to traditional presets in photo editing software, effectively addressing issues of visual inconsistency.
2 Related Work
Style Transfer for Image Editing
Style transfer was first proposed [40, 15] as an idea to transfer the style of a reference image to another image. The early research was more focused on the artistic style transfer [7, 13, 56] that altered both the textures and colors of the input images. These networks were developed for research and artistic purposes, thus, they do not have fidelity constraints i.e. the methods can alter substantial structural attributes of the scene, add new elements, change notably the colors, etc.
Photo-realistic style transfer focuses on applications where we need to maintain high-fidelity w.r.t. the input image [32, 54, 3]. These methods constrain the model by using strong regularization e.g. pixel-wise operations and losses.
The most related work focuses on color transfer [40, 39, 53, 19, 23], which limits the style transfer to the overall colors of the reference image. These models do not change the (structural) content of the input image, but they are mostly limited to global color and tone modifications.
We can highlight methods based on 3D LUT for global tone mapping and color manipulation [55, 52, 30, 48], and other similar methods such as Deep Preset [19] and Neural Preset [23].
Implicit Neural Representations
Implicit neural representation (INR) [42, 16, 36] is the concept of representing information using a neural network. This approach mainly uses MLPs as the neural representation. It is widely used in computer vision and image processing applications like 3D reconstruction [16, 34, 35, 36], Image Compression [42, 44], Video Compression [8, 33], Gamut Mapping [26], Raw Reconstruction [27] and much more [11, 4, 51].
Because of the power of INR as a function estimation, it can interpolate missing information. For that reason, it is used for arbitrary scale applications in super-resolution [10], image generation [43], and optical flow [21]. Current research tries to find new INR methods for a better neural representation [42, 46, 41]. Other works try to develop general INRs that can represent multiple data with a single representation [47, 9, 24].
Example-Based Learning
Example-based learning is concerned with models that take an example that represents the required task and apply this task to an input image. Image Analogies [18, 29, 45] is a type of algorithms that utilize a pair of images that specify a transaction to be applied on an input image. Other research reframes it as a visual prompting [5, 38] and in-context learning [49, 50, 17, 37] to create a general-purpose model that extended to more tasks like segmentation, image enhancement, and style transfer. These methods utilize a pair of images as an example to learn the underlying task and generalize to new unseen ones. In this paper, we reframed the task of reference-based image editing as an Example-based method by transferring edits instead of style.
Related Datasets
Considering the importance of image enhancement and style transfer in the computational photography and computer vision community, we analyzed the existing datasets and we found multiple issues. The image dataset proposed by Neural Preset [23] is not publicly available. The 3D LUTs used to create the dataset are not available either. Additionally, this dataset is limited to color modifications applied through 3D LUTs. The dataset proposed by Deep Preset [19] is not publicly available. Moreover, the dataset is limited by the used image variety and the number of modifications included. The well-known MIT5K dataset [6] only includes 5 different photographer styles. Moreover, most of the transformations are global – many times a single 3D LUT.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input Photographer 1 Photographer 2 Photographer 3 Photographer 4 Photographer 5
Figure 2: Samples from our dataset. We can appreciate the diversity of styles and the challenging modifications.
3 The Neural Retouch Dataset
Photography edition is a time-consuming process that requires extensive knowledge of image processing and photography e.g. white balance, contrast, saturation, tones, light conditions, etc. For these reasons, professional image editing software is not accessible to the general public e.g. smartphone users. To speed up the “retouching process”, professional photographers often develop “presets”, a series of pre-defined modifications that are applied to the image to enhance it. Presets are designed to transfer edits between images. They are often applied directly on RAW images, as they have more information than the processed RGB (JPEG).
It is worth noting that the modifications within a preset can be local and global, for example, apply certain color corrections to the sky, a different white balance to dark areas, and add vignetting and fine grain. This is thanks to the integration of automatic segmentation masks in the preset.
Considering this, we aim to create a challenging and realistic dataset for image retouching and enhancement. To create our dataset we used Adobe Lightroom software. We chose 170 different presets, licensed under Creative Commons (CC 4.0). We carefully selected the presets to include a wide variety of styles and edits. We also made sure to avoid specific purpose presets (e.g. portrait presets) and geometrical edits (e.g. Cropping, Rotation).
We selected 570 RAW images from MIT5K dataset [6]. We used the dataset metadata to include diverse images with different content, locations, lighting conditions, and cameras. Using RAW images is crucial since presets (and 3D LUTs) are designed to process RAW images or images in other color spaces different than sRGB (8 bits). This was not considered in previous works [23, 19], where the preset or 3D LUT was applied directly to the RGB image. To the best of our knowledge, this is the first (open-source) dataset that includes this variety of high-quality images and styles resulting in 100.000 retouched images.
Dataset Samples
We can see in Figure 2 some of the presets that were included in our dataset. We can appreciate the variety of edits, for instance, the fine grain, spatial edits around lights, and different vignetting effects. We can also appreciate local transformations, for instance, the sky in row 2 is modified in a different manner w.r.t. to the rest of the scene.
Discussion
We noticed that the same presets can affect images differently. As we can see in Figure 2, even though we applied the same preset (e.g. Photographer 2, 5), different images are modified in very different manners. The output of the editing is highly dependent on the look of the image before editing. In the normal color style transfer methods and datasets, the general assumption is to apply edits on neutral-looking images which limits the capabilities to work with different inputs with different looks and color distributions.
During our early experiments we noticed that, due to the different behavior of the presets depending on the image, it is quite ambiguous from the model perspective (and optimization-wise) to transfer the style given just a single reference. Having the reference before and after editing highly overcomes the issue of editing ambiguity for the model and gives the user control over the edits to be applied to the input image. Additionally, it is important to develop a method that will be robust to these differences and less dependent on data consistency.
4 Methodology
Figure 3: Our proposed INR Pipeline in comparison with the MLP INR Pipeline.
In photo editing software image editing is done by changing the colors of the image based on their values, location, context, and content. Even though editing mostly changes pixel colors, these changes are more complicated than simple color LUTs or color modification matrices. We utilize Implicit Neural representation (INR) to overcome the limitations of neural network methods. Our method does not require pre-training, and it is able to learn color changes adaptively from a single reference.
Moreover, compared with previous works that use MLP-based architectures, we propose a novel INR architecture with spatial and context awareness allowing for more accurate and adaptive edits.
General Pipeline
Our reference is represented as an edited image Before $\mathcal{B}$ and After editing $\mathcal{A}$ We use the reference to train an INR that learns the edits applied to the colors based on their content, location, and context. After training we run the input image $\mathcal{I}$ through the trained INR to obtain our output $\mathcal{I_{A}}$ Because of the nature of our dataset, we have a ground truth (GT) $\mathcal{S}$ that allows us to measure the ability of the model to extract edits from the reference and apply them to new images.
Window Sampling
Let’s consider $R$ and $P$ as the RGB and its coordinates respectively. Commonly INR architecture is constructed of MLP layers. These layers process every single pixel separately so when we work with images we have to disassemble the image into individual pixels ($N=W\times H$) and sample from them $Inp=\{r_{i}^{3}\}_{i=0}^{N}$. For our task, the position of the processed pixels is important so the coordinates of the pixel are also included as part of the input $Inp^{N\times c}=\{p_{i}^{2},r_{i}^{3}\}_{i=0}^{N}$.
When dealing with images, MLP layers can be represented as convolution layers with $1\times 1$ kernels. This allows us to do operations on single pixels without disassembling the image to individual pixels and allows us to apply spatial operations on the image as part of the architecture. This design requires processing the full image to apply weight update which increases the time required for training. To overcome this issue we introduce a window sampling scheme by replacing the pixel sampling with a window sampling. Instead of sampling by choosing random pixels, we include the neighboring pixels of the sampled pixel to construct a window of size ($n\times n$) $Inp^{N\times c\times n\times n}=\{p_{i}^{2\times n\times n},r_{i}^{3\times n\times n}\}_{i=0}^{N}$, i is the center of the $n\times n$ Window. As we see in Fig. 3, we sample the windows and treat them as small image patches. We apply the same sampling process to obtain the input coordinates and the GT samples for loss computation. This allows us to have the flexibility of pixel sampling which allows weight update after processing only parts of the image allowing for a faster convergence.
Split Processing
In our task, the location of the pixel is required for some transformations such as vignetting. We include the 2D positional encoding of the pixel, additionally to its RGB value, as an input into our INR to ensure certain spatial awareness. Inspired by [27] we split the processing of the different inputs as they differ in their importance.
Context Aware Processing
Even though the location of the pixels gives some spatial awareness to the INR, the network still does not have information about the context of the pixel w.r.t. the neighboring pixels. We aim to bring locality into the INR that otherwise would process each pixel independently i.e. a pixel-wise convolution. To achieve this, we included a depth-wise convolution with $3\times 3$ kernels to give our network context awareness. We choose this layer specifically to keep the efficiency of the INR architecture and introduce as less parameters as possible to allow for fast optimization and avoid network overfitting. The full pipeline can be summarized in the next equations:
$F_{r}=Conv_{1\times 1}(S_{r}^{d,3,n,n}),F_{p}=Conv_{1\times 1}(S_{p}^{d,2,n,n})$ (1)
$F_{concat}=Concat(F_{p},F_{r})$ (2)
$F_{context}=DWConv_{3\times 3}(F_{concat})$ (3)
$Out=Conv_{1\times 1}(F_{context})$ (4)
5 Experiments
In this section, we will describe our experiment setup with some ablation studies for our INR architecture. We provide a comprehensive benchmark comparing our proposed method to different generative and editing networks adapted for our task.
Datasets
For the training set, we used 510 images with 150 different presets. For the validation set, we used the remaining 60 images with the same 150 presets from the training set. For the test set, we used a new unseen 22 presets with the images from the validation set. For our proposed method, we require no pre-training so we only used the testing setup.
Full Reference Evaluation
As mentioned in the dataset section 3, the same preset can generate a different-looking output when applied to different images. To make sure that the difference between the input and the GT visually matching to the difference in the reference pair, for each input image we choose the reference that closely matches the color distribution of the input. We do so by comparing the 3D color histogram between the input image and the reference before edit. For a fair comparison, we use the same reference for all the tested models. We use PSNR and SSIM to evaluate the accuracy of the retouch transfer.
Implementation Details
We train our INR architecture for 1000 iterations with a sampling window size of 13 and 484 samples per iteration, using the L1 loss function. We use Adam [25] optimizer with a learning rate of $1e^{-3}$ that is gradually decreased to $1e^{-4}$ using Cosine Annealing [31] learning rate scheduler. We provide more technical details related to data pre-processing, training, and reproducibility in the supplementary material.
5.1 Context-Aware INR Ablation
Table 1: Ablation study of our modifications to the traditional INR.
Method PSNR $\uparrow$ SSIM $\uparrow$
Traditional INRs MLP 22.7422 0.7957
MLP + Enc[46] 21.5339 0.7537
MLP + Residual 22.9090 0.7966
SIREN[42] 22.8655 0.7929
INRetouch (ours) SIREN + Residual 23.0531 0.7996
+ Split 23.1025 0.7998
+ Context-Awareness 23.4216 0.8054
In our experiments, we tested different kinds of INR architecture to see what would better fit our task. As seen in Tab. 1 from the traditional INRs, Siren non-linearity [42] and residual update perform the best. Siren was able to compress the reference well while being able to generalize to new images with different shapes and content.
For our proposed architecture processing color and coordinates separately proves effective when giving more attention to the pixel color. Lastly, our context awareness through the Depth-Wise CNN layer achieves the biggest improvement proving the importance of neighboring information in our task.
Figure 4: Ablation Results for Number of References and Convergence Speed
Number of References
As we see in Fig. 4 (a) Increasing the number of references improves the performance which shows the ability of our method to interpolate and extract information from multiple sources. This is effective when working with predefined styles with multiple references available.
Convergence Speed
Our window sampling technique proves effective as we see in Fig. 4 (b) achieving a similar optimization speed as pixel sampling as it provides more flexibility with the number of updates and with the additional benefits of context awareness in improving the performance. We can also notice the issue of full image training as it requires more time to optimize.
5.2 Photo Retouching Benchmark
Table 2: Performance of different models in Photo Retouching
Type Method PSNR $\uparrow$ SSIM $\uparrow$
Pretraining StyleGan [22] 20.6370 0.7587
Deep Preset [19] 21.9494 0.7727
Neural Preset [23] 22.1291 0.7602
No Pretraining Image Analogies [18] 12.3230 0.4031
Deep Image Analogies [29] 12.7634 0.3195
Painter [49] 12.2027 0.3500
Visual Prompting [5] 14.6115 0.4129
Ours 23.4216 0.8054
For a comprehensive benchmark for our task, we tested different neural network architectures, that were proposed for generative and style transfer tasks. For an accurate comparison, we adapted these networks to work with our proposed task and were trained on our dataset. Additionally, we included other example-based methods that require no retraining and were developed for general-purpose example-based applications. We tested 2 kinds of these approaches including image analogies [18, 29] and In-Context learning [49, 5] approaches. It is important to highlight all the methods compared have the same input information.
As we see from the results in Tab. 2 we achieve the best performance with a big margin over the other methods. Methods that require pertaining produce big discrepancies in performance between seen and unseen styles as they struggle to generalize for new unseen styles.
The methods that require no pre-training struggle as they are not robust enough to learn the task from an example, even though these methods are developed for example-based learning. In the image analogies method, they work by copying from the reference. This limits their abilities to the information in the reference without the ability to generalize or interpolate any missing information. The In-Context learning methods were developed to generalize to unseen tasks but they fail to generalize for our task as they are limited to tasks very closely related to their training data. Even though Visual Prompting [5] has style transfer and Painter [49] has low light image enhancement in these tasks they still fail to recognize the required task.
Our proposed method performs the best requiring no pre-training which allows it to generalize to edits and styles without any issue. Additionally, it learns the task required from only the given example which avoids the task confusion present in the other In Context learning approaches. Additionally, our method is by far the most efficient with only 11.5 K parameters which makes it very practical and allows for high-resolution image editing in very limited hardware. We provide an extensive efficiency study of our method in the supplementary materials.
5.3 Context-aware INR on different tasks
To show the importance of context awareness and the effectiveness of our proposed architecture in image reconstruction, we test our architecture on a variety of image reconstruction applications that utilize INRs. We test our architecture on Gamut Mapping [26], Metadata-Based RAW Reconstruction [27], Neural Implicit LUT [11].
Table 3: Results of adding our proposed context-awareness to INR-based image reconstruction tasks.
Type Method PSNR $\uparrow$ SSIM $\uparrow$
Gamut Mapping[26] SIREN 54.4262 0.9997
+ Context-Awareness 55.8516 0.9998
RAW Reconstruction[27] SIRE 48.3956 0.9965
+ Context-Awareness 49.1699 0.9963
Neural LUT[11] SIREN 33.9761 0.9632
+ Context-Awareness 35.1374 0.9681
We show the difference between traditional INR architecture with pixel sampling and MLP layers, in comparison with our method that employs window sampling and context awareness. The difference between the 2 tested architectures is the sampling technique and the addition of context awareness. We did not include split processing in this comparison as we wanted to focus on the importance of context in these tasks. As we see in Tab. 3, our proposed INR architecture consistently performs better, showing the importance of context for image-related tasks. Our method proves effective in different tasks while maintaining the advantages of INR for efficiency and speed.
5.4 Qualitative Results
Figure 5: Editing consistency in our method and Lightroom preset.
Figure 6: Comparison between different INR architectures for our retouching transfer task
INR architecture comparison
As we see in Fig. C, we can notice the superiority of our proposed INR architectures. INR architecture using Fourier Feature encoding as input produces a noticeable artifact as it overfits the reference data failing to generalize to other images. The other compared methods struggle with the smoothness of the editing as well as recognizing the context for more accurate color changes. As we notice in the first row our method produces smoother edits (e.g. faces) and edits colors differently based on the context and content (e.g. chair). In the second row, we can notice the same effect by recognizing the context and content (e.g. trees, sky).
Figure 7: Comparison between different methods for our retouching transfer task
Comparison with other methods
In Fig. 7 we show the quality of our output compared to other methods. We can appreciate the consistency of our method with the ground truth producing a high-quality output without any artifacts. We can notice the ability of our method to learn complex edits accurately like the smooth vignetting effect (row 1) and blurring effect (row 3). Because our method does not require pre-training and learns directly from the reference, it has no limitation on the edition ability. It can produce local and content-based edits accurately (rows 2 and 4). On the contrary, the methods that require pre-training are limited to the training dataset and might fail to generalize to new edits (row 3). Additionally, we can notice generated artifacts in their outputs failing to apply smooth edits and high-quality output (e.g. vignetting row 1, sky row 2).
In-context learning methods such as Painter [49], fail to generalize on our task because this requires output fidelity without producing noticeable artifacts. Additionally, Image Analogies [29] are limited to the reference information producing unreliable output. Our proposed method is able to overcome the limitations of previous methods.
Figure 8: Comparison between using a style reference, and a Before-After edited reference.
Comparison with one-shot
The current style transfer method uses a single reference image to represent the style of the desired output. As we see in Fig. 8, these methods fail in the task of photo retouching. The task of photo retouching requires fine edits and specific color changes based on location and context. It is not feasible to capture these edits using only a single reference image. We tested different style transfer methods developed for photo editing based on a reference. We can notice artifacts on the output produced by undesirable changes. Additionally, they fail to recognize the fine details of the style limited to reference ambiguity. For a quality output, We notice these methods are limited to reference images with similar characteristics to the input image (nature, portrait) or with general and noticeable aesthetics (color filter, day-night images, etc). Our proposed approach allows the use of any available reference with much less limitation for high-quality output.
Visual Style Consistency
As mentioned in Section 3 the biggest challenge working with the proposed dataset is the inconsistency with the preset output, which complicates training. This inconsistency is also an issue of presets in photo editing software for the end user. They provide inconsistent visual edits when using the same preset across different images, which limits their editing transfer capabilities. Our method can serve as an alternative for edit transfer providing more consistent visual edits – see Fig. B. It can generalize to new edits without any changes or retraining required as it learns adaptively from the given reference. Additionally, it will provide a platform-independent method for edit transfer between different editing software. We discuss this into details in the supplementary material.
6 Conclusion
We propose a novel image enhancement problem focused on image editing and retouching images. First, we present a novel dataset for image retouching that poses new challenges and improves previous datasets such as MIT-5K. Second, we propose a novel method that employs context-aware implicit neural representations (INRs) for learning complex image editions. Our results show the potential of our approach as a general image retouching method, improving the performance of INR on different image restoration tasks. Moreover, generative methods for image editing could also benefit from our dataset and task. Our code and datasets will be open-source upon acceptance.
Looking forward, this work opens new possibilities for research in adaptive image manipulation, suggesting that complex editing operations can be effectively learned and transferred without sacrificing quality or control. We believe our contribution provides a foundation for future work in automated photo editing.
[…]
Supplementary Material
Method # Params (M) HD Full-HD 2K 4K # PSNR (dB)
MACs (G) Time (s) MACs (G) Time (s) MACs (G) Time (s) MACs (G) Time (s)
Deep Preset[19] 36.6 512.14 0.0578 1160.84 0.1542 2048.53 0.2820 4609.19 0.6601 21.9494
StyleGan [22] .61524 24.86 0.0316 55.95 0.0681 99.4 0.1380 223.78 0.2926 20.6370
Neural Preset [23] 4.8 153.87 0.0902 348.77 0.2355 615.47 0.4293 OOM OOM 22.1291
SIERN [42] 0.00800 8.2 0.0055 18.45 0.0124 32.81 0.0191 73.81 0.0459 22.8655
SIERN-Split [27] 0.01085 10.0 0.0087 22.5 0.0219 40.0 0.0301 90.0 0.07617 23.1025
INRetouch (Ours) 0.01149 10.59 0.0089 23.83 0.0215 42.36 0.0355 94.72 0.0759 23.4216
Table A: Efficiency study. We compare the models’ complexity in terms of parameters, operations (MACs), and inference time on HD (1280×720), Full-HD (1920×1080), 2K (2560×1440), and 4K (3840×2160) images. We measured inference time on the NVIDIA RTX 4090 24GB GPU. “OOM” means out-of-memory issue. The units used “s”, “G”, and “M” are seconds, gigabytes, and millions, respectively.
We first kindly refer the readers to our accompanying video, with an explanation in Sec. 1.
Then, in this supplementary material, we provide more implementation details of our work in Sec. 2. We also provide more ablation studies in Sec. 3. The efficiency study can be found in Sec. 4
As for visuals, we first provide a comparison on visual consistency in Sec. 5. More visuals on retouching transfer comparison can be found in Sec. 6. Finally, in Sec. 7, we show the variety of our presets applied to a natural image.
1 Accompanying Video
We train our model on image pairs before and after editing. Designed to be lightweight and efficient, our model enables affordable inference, which motivated us to extend its application to video editing. As shown in the accompanying video, our method effectively learns edits from images and applies them to videos, producing visually pleasing results with excellent temporal consistency and no noticeable artifacts. This can be attributed to the clarity from the use of image pairs before and after editing as a reference and the design of our method to focus on color modification through local awareness.
Unlike existing methods, such as style transfer and generative-based models, which often struggle with temporal consistency and introduce significant noise, our approach overcomes these limitations. This demonstrates both the effectiveness of our network and the controllability of the learned edits.
2 Implantation Details
Compared Methods
For the compared method that requires pre-training on the dataset, we modified and adapted their architectures for our task. For the Deep Preset [19] method, We modified the reference branch to take 6-channel input. We provide the image pair before and after editing as a reference by stacking them together. For Neural Preset we modified the architecture to generate an editing mask with the same size of the input instead of just a modification vector to allow for local modification. Similarly, we use the pair of before and after editing stacked together as the reference to the model. For the Style GAN [22] based method we used the Domain Alignment Module proposed in [14]. This module was proposed to apply color changes to an image based on a provided reference. We modified the module to take the stacked pair of before and after editing as a reference. We emphasize that all the compared methods take the same input information (reference before and after editing).
For the other methods that require no pre-training on our dataset (Image Analogies [18, 29] and In-Context learning [49, 5] methods), we used the open-source models provided by the authors.
Evaluation Dataset
Lightroom preset system suffers from visual inconsistency. As we see in Fig. B same preset can produce different styles when applied to different images. For an accurate evaluation process, we need to make sure the chosen reference visual style matches the style of the GT. We achieve that by choosing a reference that has the same color distribution as the input image as it is more likely to generate the same style when applying the preset. We calculate the 3D color histogram of each reference image before editing and we compare it with the 3D color histogram of the input image. We choose the reference image with the closest color histogram to the input images as a reference. For a fair comparison, We used the same reference in all compared methods.
Dataset Disclaimer
All the images used in our dataset were obtained from the open-source MIT5k dataset [6] and all the presets used are open-access licensed under Creative Commons (CC 4.0). All dataset creation processes and components are checked to avoid any violations or misuse and to ensure ethical conduct.
3 More Ablations
Figure A: Ablation Results for Sampling Window Size and DW CNN Filter Size.
Sampling Window Size
In Fig. A (a) we can notice some improvement when increasing the size of the sampled window. This can be attributed to the model processing a bigger coherent area to learn more about update smoothness. But after some degree, we see no noticeable improvement. For our experiments, we chose a sampling Window size of 13 for the best trade-off between cohesion and memory footprint during training.
DW CNN Filter Size
Fig. A (b) shows increasing filter size can improve performance as it considers more information from neighboring pixels. However, increasing the filter size introduces more parameters that require more time to optimize and can result in overfitting issues with a drop in performance. We choose the filter size of 3 for fast optimization and as less parameters increase as possible.
4 Efficiency Study
As we see in Tab. A, our proposed method is very efficient with very few parameters (11k) requiring very little memory and with a fast inference time, processing a 4K image in just 70 ms. The style transfer methods compared (first 3 rows), have much more parameters (60x to 400x more parameters) and need 5x to 10x more time to process the same image. Additionally, in comparison to the INR methods we can notice our model achieving a similar efficiency and inference speed. Even though our new INR method requires more parameters, our model design and layer choices were able to maintain a similar efficiency while noticeably improving the performance.
5 Visual Style Consistency
Figure B: Editing consistency in our method and Lightroom preset.
In Fig. B, we compare the editing consistency of our proposed method with that of the widely-used Lightroom preset. Our method demonstrates the ability to produce more realistic outputs that better adhere to the reference edits, validating its effectiveness and highlighting its superior visual consistency.
Lightroom presets work by saving the Lightroom edits applied to an image. These presets are usually created to process RAW images. These edits consist of image processing pipeline operations like color correction, hue adjustment, and exposure correction. These operations affect every image differently depending on the image details like the sensor of the camera, lighting conditions, and color distribution. This limits the visual reproducibility of the edits to similar images. Additionally saving edits in formats like presets is software-specific so they require the same software to use them. Our method provides a more visually consistent way to transfer edits between images without being software-dependent.
6 More Visual Results
We show in Fig. C the qualitative results of various methods for the retouching transfer task. Our approach excels in accurately learning the edits from before-and-after image pairs, producing outputs that are not only more realistic but also better aligned with the intended edits. In contrast, other methods struggle to achieve similar fidelity, often resulting in noticeable artifacts and inconsistencies. This highlights the effectiveness and robustness of our method in capturing and applying complex retouching transformations.
Figure C: Comparison between different methods for our retouching transfer task
7 Presets for Our Dataset
To ensure the versatility and robustness of our dataset, we curated a diverse collection of varying presets, designed to simulate a wide range of editing styles and conditions. As shown in Fig. D, we apply some of these presets to a single natural image, showcasing the richness and variety inherent in the dataset. This comprehensive coverage not only highlights the adaptability of our approach to diverse editing scenarios but also establishes our dataset as a valuable resource for developing and evaluating methods capable of handling complex retouching tasks. Such diversity enables the models trained on our dataset to generalize effectively across different styles. We showcase the ability of our method to simulate these presets on videos in our accompanying video.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure D: Visualization of the output of our different presets applied to a natural image (highlighted top left). Zoom in to see better.